OpenStudy (anonymous):

Algebra Questions- I really need help.... 2. Simplify the radical expression square root 72x^2 A) 36 square root x^2 B) 6x square root 2 C) x square root 36 D) 6 square root 2x^2 3. Simplify the radical expression. square root 150x^3k^4 A) 75xk^2 B) 5xk square root 6xk C) 5xk^2 square root 6x D) 5 square root 6x^3k^4

OpenStudy (anonymous):

For 2:\[\sqrt{72x ^{2}}\] Can you put 72 into prime factor form? As in, break it down into multiplication of all prime numbers.

OpenStudy (anonymous):

@Alphabet_Sam Factors of 72: 1,2,3,4,6,8,9,12,18,24,36,72

OpenStudy (anonymous):

\[\sqrt{72}=\sqrt{8}\sqrt{9}\]

OpenStudy (anonymous):

\[6x \sqrt{2}\]

OpenStudy (anonymous):

Square root 8: 2.8 Square root 9: 3 @SkepticGod

OpenStudy (anonymous):

you cant squartroot 8 because then decimal so keep it like above i did

OpenStudy (anonymous):

@SkepticGod Oh okay, can you help me with a few more please

jhonyy9 (jhonyy9):

2. sqrt72x^2 = sqrt(36*2*x^2) = sqrt(6^2*2*x^2) = ? the perfect squares take out from squarroot so than what will get ?

OpenStudy (anonymous):

sure so for problem#3 the same concept. break down large number into smaller number that you could squareroot without getting decimal

jhonyy9 (jhonyy9):

@Answers101 so do you understand it why are in this way ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\sqrt{150}=\sqrt{25}\sqrt{6}\]

OpenStudy (anonymous):

since x is cube you can divide like this \[\sqrt{x^3}=\sqrt{x^2}\sqrt{x}\]

OpenStudy (anonymous):

same thing for k^4 \[\sqrt{k^4}=\sqrt{k^2}\sqrt{k^2}\]

OpenStudy (anonymous):

@Answers101 what do you think the answer is?

OpenStudy (anonymous):

ops nevermind too late typing